# Characterizing GPU Resilience and Impact on AI/HPC Systems

### v3

@@ 4.3. Examples of Error Propagation to User Jobs
Incident 1: A GSP error stalled GPU control functions and rendered the GPU inoperable on an A100 GPU. Consequently, the user job scheduled on that GPU failed. The GSP error required the draining of the node and a full node reboot to recover, which led to the draining of all pending user jobs on that node. From the beginning of the node drain to the completion of the node reboot, the total recovery time for this incident was 23 node hours (09:08 AM to 08:21 AM the next day), during which the node was unavailable. Figure 9(a) shows that the corresponding GPU utilization dropped quickly following the error incident due to job failure. This incident shows that a GPU error can significantly interrupt user jobs and node availability.
Incident 2: In this incident, the H100 GPU running the user job experienced an uncorrectable ECC memory error, which led to an error containment event. The error containment event contains the uncorrectable memory error by terminating the user’s job. Subsequently, a node reboot is triggered to recover from this error, resulting in a two minute node downtime. Figure 9(b) shows that GPU utilization dropped as the user job was terminated. This incident shows that although an uncorrectable memory error might be contained, it still results in user job termination and requires node reboot to fully recover, resulting in unexpected node downtime.
@@ A GPU error can lead to:
  1. (1)
    Job Failure: A GPU error may not be handled by jobs, either because the error itself is not contained or because there is a lack of appropriate error-handling mechanisms. For example, GSP errors lead to GPU failures and require node reboots. To recover from GPU failures, jobs need to be re-executed from the beginning or rolled back to the closest checkpoint.
  2. (2)
    GPU and Node Downtime: Downtime can occur when GPUs need to be reset or replaced by the operator, and no jobs can be scheduled on the GPU and the corresponding node in the interim.
@@ 5.1. Result Highlights
(i) Except for MMU and NVLink errors, no other GPU errors are handled by jobs, thus resulting in their failure. Depending on the GPU type, the underlying cause of job failures differs. In the case of A100 GPUs, hardware errors predominantly lead to job failures, whereas memory errors are the primary cause of H100 GPUs. Therefore, there is a need to improve the resilience of underlying hardware to minimize such failures. While checkpointing is an option, checkpointing routines have a high overhead of up to 40% (Wang et al., 2005; Wan et al., 2024; Maurya et al., 2024), including management, storage, and restore.
(ii) The overall availability per-GPU node is $\sim$99.4(corresponding to 9 minutes downtime/day) and $\sim$99.3(corresponding to 10 minutes downtime/day) for A100 and H100 GPUs respectively. This level of unavailability suggests that even if the rest of the infrastructure is highly available, current GPUs may not provide sufficient availability to meet the demands of critical applications that require greater than 3 9’s of availability (downtime of 1.4 mins/day). In addition, we used emulation to project the impact of this availability distribution at increased scales (for both node scale and job duration) and found that significant overprovisioning of 5% would be necessary to handle associated failures (as explained in further detail in Section 5.4).
@@ 5.3. Job Failure Analysis
XID GPU Error # GPU-failed Jobs # Jobs Encountering Given XID Job Failure Probability (%) A100 H100 A100 H100 A100 H100 31 MMU err. 3206 93 3543 126 90.48 73.80 74 NVL err. 43 0 80 0 53.75 122 SPI PMU RPC failure 40 0 41 0 97.56 119 GSP RPC timeout 31 0 31 0 100.00 94 Contained ECC 5 5 5 5 100.00 100.00 48 GPU DBE 0 5 0 5 100.00 64 Row remapping failed 0 3 0 3 100.00 63 Row remapping event 0 2 0 2 100.00
@@
NVLink and MMU errors do not necessarily lead to job failures because:
  1. (1)
    For NVLink errors, the link or GPU may not be in use by any user jobs (as discussed in Section 4.2) [^12] Based on our understanding, NVLink and memory errors can occur even when no workload is running (Forums, 2024)..
  2. (2)
    For MMU errors, there can be application or library-level masking mechanisms. Besides hardware errors, MMU errors can also occur if buggy user code makes illegal memory accesses that cannot be mapped in the virtual-to-physical address space. Such errors can be managed using appropriate application-level exception handlers. Popular libraries and frameworks for machine learning (PyTorch Contributors, 2018, 2017; PyTorch Lightning Contributors, 2021) have support for handling such exceptions by skipping the associated training iteration, albeit at the cost of model quality.
@@ 5.4. Impact of GPU Downtime on Jobs
The simulation uses a discrete time event simulation with node failure probabilities derived from our prior analysis. The recovery time after a failure is dependent on variables such as checkpoint load time and availability of spare GPUs. To account for the variability introduced by these factors, we parameterize recovery time and perform a parameter sweep. For a training job with 608 GPUs and recovery time of 2.2 hours, the required overprovisioning is 5%: i.e., 31 additional GPUs are needed beyond the original 608 to maintain availability of 99.9% at the job level. While at first glance, such overprovisioning would appear to be a small cost, for the above example, it would cost over \$1 million per month for a 1000 node cluster (our analysis is based on AWS H100 GPU rental rates (DataCrunch, 2024)). However, if the recovery time is reduced to 5 minutes, downtime decreases significantly, and the required overprovisioning drops to 2%, a 2.5×2.5$\times$2.5 × reduction. This highlights the criticality of minimizing recovery time to reduce downtime for large and long-running jobs.
@@ 6. Related Work
Existing work has analyzed GPU resilience at the microarchitecture, system, and application levels. This paper extends previous work via comprehensive analyses of GPU error characteristics, failure propagation, and their impact on user jobs.
@@
Microarchitecture-level GPU Resilience. Previous research (Vallero et al., 2018; Sartzetakis et al., 2022; Hari et al., 2015; Yang et al., 2024) has primarily focused on the resilience of individual GPUs at the microarchitecture and software levels, e.g., for older generations such as the NVIDIA G80 (Braga et al., 2023). However, the earlier work did not evaluate the resiliency of modern GPUs in large-scale HPC settings.
@@
System-level GPU Resilience in HPC Settings. Existing studies have analyzed the resilience of GPUs in HPC systems (Haque and Pande, 2010; DeBardeleben et al., 2014; Kokolis et al., 2025; Gunawi et al., 2018; Cini and Yalcin, 2020), for example, the NVIDIA Tesla K20X GPUs in various supercomputers (Di Martino et al., 2014; Tiwari et al., 2015b, a; Gupta et al., 2015; Nie et al., 2016; Gupta et al., 2017; Nie et al., 2017; Ostrouchov et al., 2020). Studies of the Blue Waters, Titan, and Summit supercomputers (Di Martino et al., 2014; Nie et al., 2016, 2017; Oles et al., 2024; Kokolis et al., 2025) have examined node failures and GPU error characteristics. However, such work either studied previous generations of GPUs with a focus on GPU memory errors, or focus on cluster-level resilience instead of GPU resilience. Our work complements those efforts by providing resilience insights into the latest generation of GPUs, focusing on a broader range of components.
@@
Application-level GPU Resilience in Data-centers and Deep Learning Workloads. Recent research has focused on understanding GPU power usage (Nie et al., 2017), GPU component-level failures (Jung and Erez, 2023), software-level error handling (Maruyama et al., 2010; Fang et al., 2014; Lian et al., 2025), and the impact of GPU software error propagation on GPGPU applications (Wei et al., 2020; Li et al., 2016; Anwer et al., 2020; Rech et al., 2014) and emerging GPU workloads such as convolutional neural networks (CNNs) (dos Santos et al., 2018; Condia et al., 2022), large language models (LLMs) (Jiang et al., 2024; Dubey et al., 2024), and safety-critical applications (Perez-Cerrolaza et al., 2022; Shah et al., 2024).
7. Discussion
Justification of Analyzing Errors: Data-driven HPC resilience characterization studies analyze operational data on system/application errors to provide insights into system resilience (DeBardeleben et al., 2014; Di Martino et al., 2014; Tiwari et al., 2015b, a; Gupta et al., 2015, 2017; Ostrouchov et al., 2020; Nie et al., 2016, 2017; Oles et al., 2024). While in those studies, the error rate is used as the key metric to quantify resilience, some (Sridharan et al., 2015; Beigi et al., 2023) argue that fault rate is a more appropriate metric.
Errors represent the manifestation of faults and have direct downstream consequences, such as triggering of recovery mechanisms, application interruptions, or system-wide outages (SWOs). While a fault may result in multiple errors, it is the resulting errors that the recovery mechanisms must address to maintain system health. These errors and their recovery process directly impact system health, performance, and availability. Thus, SREs prioritize errors over faults. Hence, like many others who study operational data, we chose to study errors.
Potential errors in NVIDIA driver logging: A potential source of error is the NVIDIA driver’s inconsistent logging of different XID errors. The NVIDIA driver is proprietary software whose source code is not publicly available, making exact specifications of its error-logging mechanism unverifiable. We attempted to minimize the impact of the inconsistency, to the best of our ability, through consistent log data collection and error coalescing.

### v4

@@ 4.3. Examples of Error Propagation to User Jobs
Incident 1: A GSP error stalled GPU control functions and rendered the GPU inoperable on an A100 GPU. Consequently, the user job scheduled on that GPU failed. The GSP error required the draining of the node and a full node reboot to recover, which led to the draining of all pending user jobs on that node. From the beginning of the node drain to the completion of the node reboot, the total recovery time for this incident was 23 node hours (09:08 AM to 08:21 AM the next day), during which the node was unavailable. Figure 9a shows that the corresponding GPU utilization dropped quickly following the error incident due to job failure. This incident shows that a GPU error can significantly interrupt user jobs and node availability.
Incident 2: In this incident, the H100 GPU running the user job experienced an uncorrectable ECC memory error, which led to an error containment event. The error containment event contains the uncorrectable memory error by terminating the user’s job. Subsequently, a node reboot is triggered to recover from this error, resulting in a two minute node downtime. Figure 9b shows that GPU utilization dropped as the user job was terminated. This incident shows that although an uncorrectable memory error might be contained, it still results in user job termination and requires node reboot to fully recover, resulting in unexpected node downtime.
@@ A GPU error can lead to:
(i) Job Failure: A GPU error may not be handled by jobs, either because the error itself is not contained or because there is a lack of appropriate error-handling mechanisms. For example, GSP errors lead to GPU failures and require node reboots. To recover from GPU failures, jobs need to be re-executed from the beginning or rolled back to the closest checkpoint.
(ii) GPU and Node Downtime: Downtime can occur when GPUs need to be reset or replaced by the operator, and no jobs can be scheduled on the GPU and the corresponding node in the interim.
@@ 5.1. Result Highlights
(i) Except for MMU and NVLink errors, no other GPU errors are handled by jobs, thus resulting in their failure. Depending on the GPU type, the underlying cause of job failures differs. In the case of A100 GPUs, hardware errors predominantly lead to job failures, whereas memory errors are the primary cause of H100 GPUs. Therefore, there is a need to improve the resilience of underlying hardware to minimize such failures. While checkpointing is an option, checkpointing routines have a high overhead of up to 40% (wang2005modeling; wan2024bytecheckpoint; maurya2024datastates), including management, storage, and restore.
(ii) The overall availability per-GPU node is $\sim$99.4% (corresponding to 9 minutes downtime/day) and $\sim$99.3% (corresponding to 10 minutes downtime/day) for A100 and H100 GPUs respectively. This level of unavailability suggests that even if the rest of the infrastructure is highly available, current GPUs may not provide sufficient availability to meet the demands of critical applications that require greater than 3 9’s of availability (downtime of 1.4 mins/day). In addition, we used emulation to project the impact of this availability distribution at increased scales (for both node scale and job duration) and found that significant overprovisioning of 5% would be necessary to handle associated failures (as explained in further detail in Section 5.4).
@@ 5.3. Job Failure Analysis
XID GPU Error # GPU-failed Jobs # Jobs Encountering Given XID Job Failure Probability (%)
A100 H100 A100 H100 A100 H100
31 MMU err. 3206 93 3543 126 90.48 73.80
74 NVL err. 43 0 80 0 53.75
122 SPI PMU RPC failure 40 0 41 0 97.56
119 GSP RPC timeout 31 0 31 0 100.00
94 Contained ECC 5 5 5 5 100.00 100.00
48 GPU DBE 0 5 0 5 100.00
64 Row remapping failed 0 3 0 3 100.00
63 Row remapping event 0 2 0 2 100.00
@@
NVLink and MMU errors do not necessarily lead to job failures because: (1) For NVLink errors, the link or GPU may not be in use by any user jobs (as discussed in Section 4.2) [^13] Based on our understanding, NVLink and memory errors can occur even when no workload is running (nvidia_forum_issue).; and (2) For MMU errors, there can be application or library-level masking mechanisms. Besides hardware errors, MMU errors can also occur if buggy user code makes illegal memory accesses that cannot be mapped in the virtual-to-physical address space. Such errors can be managed using appropriate application-level exception handlers. Popular libraries and frameworks for machine learning (pytorch_distributed_exception_handling; pytorch_issue_1137; pytorch_lightning_discussion_15188) have support for handling such exceptions by skipping the associated training iteration, albeit at the cost of model quality.
@@ 5.4. Impact of GPU Downtime on Jobs
The simulation uses a discrete time event simulation with node failure probabilities derived from our prior analysis. The recovery time after a failure is dependent on variables such as checkpoint load time and availability of spare GPUs. To account for the variability introduced by these factors, we parameterize recovery time and perform a parameter sweep. For a training job with 608 GPUs and recovery time of 2.2 hours, the required overprovisioning is 5%: i.e., 31 additional GPUs are needed beyond the original 608 to maintain availability of 99.9% at the job level. While at first glance, such overprovisioning would appear to be a small cost, for the above example, it would cost over \$1 million per month for a 1000 node cluster (our analysis is based on AWS H100 GPU rental rates (gpu_costs)). However, if the recovery time is reduced to 5 minutes, downtime decreases significantly, and the required overprovisioning drops to 2%, a $2.5\times$ reduction. This highlights the criticality of minimizing recovery time to reduce downtime for large and long-running jobs.
@@
6. Discussion
Justification of Analyzing Errors. Data-driven HPC resilience characterization studies analyze operational data on system/application errors to provide insights into system resilience (debardeleben2014gpu; di2014lessons; tiwari2015understanding; tiwari2015reliability; gupta2015understanding; gupta2017failures; ostrouchov2020gpu; nie2016large; nie2017characterizing; oles2024understanding). While in those studies, the error rate is used as the key metric to quantify resilience, some (vilas_2015; Beigi2023) argue that fault rate is a more appropriate metric. Errors represent the manifestation of faults and have direct downstream consequences, such as triggering of recovery mechanisms, application interruptions, or system-wide outages (SWOs). While a fault may result in multiple errors, it is the resulting errors that the recovery mechanisms must address to maintain system health. These errors and their recovery process directly impact system health, performance, and availability. Thus, SREs prioritize errors over faults. Hence, like many others who study operational data, we chose to study errors.
Reliability of Logging. A potential source of error is logging inconsistency from (i) missing logs due to storage failures, (ii) node lockups where the NVIDIA driver fails to log the error prior to the failure, and (iii) incorrect job-status captured by Slurm. However, these issues negligibly affected our analysis because: (1) Delta minimizes storage failure impact by streaming logs in real time to centralized storage; (2) SRE records show only 27 A100 and nine H100 node lockups–0.18% and 0.49% of total GPU errors–making their impact negligible; and (3) while eliminating false-negative job status is challenging, GPU errors studied led to job failure almost 100% of times (Table 2), indicating failure statuses were reliably captured by Slurm and false-negatives minimally impacted our findings.
@@ 7. Related Work
Existing work has analyzed GPU resilience at the microarchitecture, cluster, and application levels. This paper extends existing work via comprehensive analyses of GPU error characteristics, propagation, and impact on user jobs.
@@
Microarchitecture-level GPU Resilience. Previous research (vallero2018multi; sartzetakis2022gpufi; hari2015sassifi; yang2024gpu) has primarily focused on the resilience of individual GPUs at the microarchitecture and software levels, e.g., for older generations such as the NVIDIA G80 (braga2023software). However, the earlier work did not evaluate the resiliency of modern GPUs in large-scale HPC settings.
@@
System-level GPU Resilience in HPC Settings. Existing studies have analyzed the resilience of GPUs in HPC systems (haque2010hard; debardeleben2014gpu; kokolis2024revisiting; gunawi2018fail; 10.1145/3372790_comment), for example, the NVIDIA Tesla K20X GPUs in various supercomputers (di2014lessons; tiwari2015understanding; tiwari2015reliability; gupta2015understanding; nie2016large; gupta2017failures; nie2017characterizing; ostrouchov2020gpu). Studies of the Blue Waters, Titan, and Summit supercomputers (di2014lessons; nie2016large; nie2017characterizing; oles2024understanding; kokolis2024revisiting) have examined node failures and GPU error characteristics. However, such work either studied previous generations of GPUs with a focus on GPU memory or on cluster-level resilience instead of GPU resilience. Our work complements those by providing resilience insights into the latest generation of GPUs, focusing on a broader range of components.
@@
Application-level GPU Resilience in Data-centers and Deep Learning Workloads. Recent research has focused on understanding GPU power usage (nie2017characterizing), GPU component-level failures (jung2023predicting), software-level error handling (maruyama2010high; fang2014gpu; lian2025), and the impact of GPU software error propagation on GPGPU applications (wei2020g; li2016understanding; anwer2020gpu; rech2014impact) and emerging GPU workloads such as convolutional neural networks (CNNs) (dos2018analyzing; condia2022multi), large language models (LLMs) (jiang2024megascale; dubey2024llama), and privacy/safety-critical applications (li2024advances; perez2022gpu; shah2024characterizing).
@@ 8. Conclusion
Acknowledgements.
We thank the reviewers for valuable feedback and J. Applequist, H. C. Fairow, S. Weick, and K. Atchley for support. This work used NCSA Delta and DeltaAI compute resources supported by NSF grants 2005572 and 2320345, with allocations from ACCESS (NSF 2138259, 2138286, 2138307, 2137603, 2138296), and additional support from NSF grants 2029049, 2319190, 2430244, 2530738, the IBM-Illinois Discovery Accelerator Institute, and the VinUni-Illinois Smart Health Center. Any opinions expressed here are those of the authors and do not reflect the views of NSF, IBM, or VinUni.